Engineers must consider the breadths of male heads when designing helmets. The company researchers have determined that the population of potential clientele have head breadths that are normally distributed with a mean of 6.7-in and a standard deviation of 0.8-in. Due to financial constraints, the helmets will be designed to fit all men except those with head breadths that are in the smallest 3.9% or largest 3.9%.

What is the minimum head breadth that will fit the clientele?
min =

What is the maximum head breadth that will fit the clientele?
max =

Enter your answer as a number accurate to 1 decimal place. Answers obtained using exact z-scores or z-scores rounded to 3 decimal places are accepted.
Answer #1

Given that,

P(Z < -1.762) = 0.039

z = -1.762

Using z-score formula,

x = z * +

x = -1.762 * 0.8 + 6.7 = 5.3

min = 5.3

P(Z < 1.762) = 0.039

z = 1.762

Using z-score formula,

x = z * +

x = 1.762 * 0.8 + 6.7 = 8.1

max = 8.1